Can you please solve this for me? I have to solve this using factoring:
The sum of the squares of two consecutive negative even integers is 340. Find the integers.

Answer by xcentaur(357)

sum of the squares of two consecutive negative even integers is 340


let the first number be x
it must be negative and even,so it becomes '-2x'
Then the second number becomes '-2x+2'


So we get,
(-2x)^2+(-2x+2)^2=340
4x^2+(4x^2+4-8x)=340
8x^2-8x+4-340=0
8x^2-8x-316=0


Solve for x. then '-2x' and '-2x+2' give you the required numbers